UT 20231129: ECD PEPSI Pol Checkout, DDT, Other


ECD was postponed until polarimeter could be installed. A full night trade on 20231129 was made with LBTB

Night Info

For the night of: Tue, 2023 Nov 28 ---> Wed, 2023 Nov 29
Local midnight = 2023 Nov 29,  7 hr UT, or JD 2460277.792
Local Mean Sidereal Time at midnight =   4 12 05.9

Sunset ( 1000 m horizon):   17 19 MST; Sunrise:   6 57 MST
Evening twilight:  18 40 MST;  LMST at evening twilight:  22 52
Morning twilight:   5 35 MST;  LMST at morning twilight:   9 48
12-degr twilight: 18 11 MST -->  6 05 MST; night center:   0 08 MST

Moonrise:  18 10 MST   
Moon at civil midnight: illuminated fraction 0.957
1.9 days after full moon, RA and dec:   6 01 56, +28 01.0

The sun is down for 13.6 hr; 10.9 hr from eve->morn 18 deg twilight.
Bright all night (moon up from evening to morning twilight).

Weather: Mostly cloudy predicted with decreasing clouds through the night

Staff

  • OSA: John Hill
  • ISA/SO: Andrew Cardwell, Justin Rupert
  • Special guests: Ilya Ilyin
  • ECD Coordinator: Jenny Power


Night Plans

PEPSI-Pol AGW Checkout (Pols not installed - unavailable)

  • AGW 3/4 Alignment
    • verify SH spots illuminated - critical - (0.5 hr)
    • PEPSI-Pol collimation check - critical - (0.5hr)
    • PEPSI POL Science Calibrator
        • Gamma Equ (RA 10hr dec 10deg) Ideal Calibrator, other options available if open later.

DDT Observations

  1. Mannucci (high priority)
    • LUCI AO on dual AGN at 3hr +41
    • Seeing <0.9" required
    • Scripts located in Service_Observing/2023B_ECD/LUCI/DDT/ OT: dual_v04.xml

*

The target is in the field of the Perseus Cluster. We intend to obtain high-resolution H and Ks images of this
object deep enough to detected the companion and measure its colors. A few NGS are present in the field.
The best one has Rp=10.34 at 47.8” of separation. The SOUL SR ratio calculator provides SR=0.15 and
FWHM=0.08” in Ks and SR=0.05 and FWHM=0.10” in H (Seeing=0.80”, zenith distance=20deg). These
figures are adequate to fully resolve the two components and measure their luminosities. Starting from the
observed optical magnitude in the VIS filter (similar to the Gaia G filter) of the fainter component,
VIS=23.35, and the typical (G-H) and (G-Ks) colors of SDSS QSO at z>1, we estimate expected
magnitudes brighter than H=21.4 and Ks=20.9 (Vega). Using the LUCI ETC, 30min per filter are enough to
obtain SNR>7 in each band assuming the SR above.

LUCI2 Transform
  1. Collect transform data - (2hrs)
  2. Target Fields, select one near transit:
    • Stone A 00 54 02 -00 32 34 Stone_A3.list
    • Stone D Ra = 04:30:27.127 Dec = -00:36:22.91 Stone_D_LUCI_2023B.list
    • Stone E Ra = 07:13:47.154 Dec = +00:13:09.40 Stone_E_LUCI_2023B.list at PA -135
    • Stone G Ra = 10:17:27.879 Dec = -01:12:47.02 Stone_G.list (short, and only to be done if shorter on time)
  3. Seeing should be no worse than ~1", list should include ~ 50 stars
  4. Have the TO send an Active preset to tune pointing on the field and collimate
  5. Configure LUCI with the N30 field stop aligned, N375 camera, H filter, flexure compensation on.
  6. Follow the instructions as outlined on: https://wiki.lbto.org/Commissioning/ComTools#Guide_Probe_Transformation
  7. Take 5xdarks to match exposure time taken
  8. Process transform data (or shoot Doug an email ASAP)
LUCI2 ESM on-sky checks
  1. AO check with LUCI2 NCPA on sky
  2. ESM Science like OB check
    1. Bin4
    2. Bin1
    3. Bin2

Closed Dome Plan

  • NCPA LUCI2 closed dome testing (high)
  • PSF Stabilization (med)
  • Generate and test low mode LBTI reconstructor for SHARK-VIS (low)

Nightlog:

PEPSI-Pol checkout done on 20231203

OSA: RA

SO: AB

Other: JH, DT, II (AIP), SJ (AIP)

00:50 Pointing check. No issues

01:00 Collimation According to John the SH spot illumination is fine. We are skipping guiding hotspot adjustment.

01:03 Preset to gamma Equ

01:10 For an unknown reason the hotspot jumps away and we lose the target on SX.

01:20 It seems like PEPSI is sending large hotspot corrections (~100px). It might be a software bug in PEPSI

01:25 Starting integrating on calibrator gamma Equ

01:40 We still have some issues with the copointing limit. Guider at the limit. John and Riccardo are manually balancing the optics while we continue observing.

02:40 Starting with LBTB science

-- JenniferPower - 27 Nov 2023
Topic revision: r6 - 03 Dec 2023, AlexanderBecker
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback